Determine the mass in grams of cycloheptane that must be added to 25 g of benzene to make a 0.62 m solution. Enter your answer to 2 decimal places.
Answer #1

mass of benzene   = 25g   = 0.025kg

molality   = no of moles of solute/weight of solvent in Kg

    0.62     = no of moles of cycloheptane /0.025

no of moles of cycloheptane   = 0.62*0.025 = 0.0155 moles of cycloheptane

mass of cycloheptane   = no of moles * gram molar mass

                                       = 0.0155*98 = 1.52g   >>>>answer